# MMformer with Adaptive Transferable Attention: Advancing Multivariate Time Series Forecasting for Environmental Applications

### v1

@@ 3 Methodology
This section is organized as follows: Section 3.1 describes the data preprocessing methods and formats. Section 3.2 comprehensively describes the methodology, including detailed procedures and clear flowcharts. Section 3.3 focuses on the model’s components, detailing the innovative components. This algorithm integrates MC Dropout regularization techniques and introduces an innovative adaptive multi-head attention mechanism. These advancements lead to substantial improvements in neural network optimization and increase the accuracy of time-series forecasting.
@@ 3.1 Data Preprocessing
  • Step 1.
    Feature sorting and selection: Features are sorted and selected using random forests to identify the prediction task’s most relevant and influential features. This step helps to reduce noise and improve the efficiency and accuracy of the model.
  • Step 2.
    Missing value and outlier processing: Missing values and outliers in the dataset are treated using techniques such as linear interpolation or MTS smoothing to ensure the continuity and reliability of the data.
  • Step 3.
    Data transformation and normalization: The selected features are subjected to necessary data transformation and normalization. This step aims to adjust the data to a form suitable for model processing.
  • Step 4.
    Integrity check: In the final stage of preprocessing, the processed data is checked for integrity, including verification of data integrity, availability, etc., to ensure that the data quality meets the requirements of model training.
@@ 3.2 Workflow Overview
This subsection outlines our research methodology, illustrated by the flow diagrams in Fig.2.
Figure 2: MMformer neural network flow chart, describing the core operating principle of the model and the operation mode of adaptive transferable multi-head attention
The MMformer uses datasets for automatic feature extraction and employs meta-learning during training to identify key sequence segments for predicting temporal patterns.
Initially, the MTS data undergoes preprocessing. Subsequently, The MMformer uses a sliding-window function to segment each city’s time series into input sequences and subsequent prediction targets, effectively creating additional subsets from the original MTS data, allowing the model to capture diverse patterns and enhance robustness and generalization.
Unlike traditional Transformers, MMformer introduces the iTransformer concept, embedding each feature’s time series into tokens to capture intervariable correlations. Our adaptive transferable multi-head attention replaces the standard self-attention mechanism, using meta-learning to initialize and quickly optimize key parameters applied to the decoder and conduct autoregressive prediction. Additionally, MC Dropout replaces conventional dropout in feed-forward networks, improving generalization, simplifying integration, and reinforcing robustness. The final phase involves an outer optimization loop that fine-tunes initial parameters through multi-task feedback, reducing overall tuning efforts.
By integrating deep learning and meta learning, the MMformer framework minimizes the need for manual feature engineering, advancing real-time forecasting. Combining insights from meteorology and machine learning, our methodology significantly improves the precision and reliability of environmental predictions.
@@ 3.3.1 Advancements in MMformer Architecture
MMformer builds upon the iTransformer framework by enhancing the self-attention mechanism and refining the network structure, as illustrated in Fig.3.
@@
Figure 4: Transformer structure
Figure 5: iTransformer structure
Compared with the traditional Transformer architecture shown in Fig.4, which relies solely on multi-head self-attention, the iTransformer (Fig.5) incorporates a "Temporal LayerNorm" layer as defined in Eq.1 [21]:
where $H$ is a matrix ($N$*$D$) of $N$ tokens, each with dimension $D$; $h_{n}$ is the sequence vector for token $n$, with $\text{Mean}(h_{n})$ and $\text{Var}(h_{n})$ representing its mean and variance, respectively. The purpose of expressing this operation in set form is to emphasize that standardization occurs separately for each time step (token). This set representation indicates that the same normalization procedure is applied independently to each time step $n$. The MMformer uses Multi-Layer Perceptrons (MLPs) for embedding and mapping layers, eliminating the need for positional encoding by independently mapping sequences to variate tokens and introducing time encoding.
MMformer enhances MTS analysis by integrating embedding methodology while retaining time positional encoding, thereby improving the processing of diverse inputs. Central to MMformer is the Adaptive Transferable Multi-Head Attention (ATMA) proposed by us, which advances traditional multi-head self-attention by dynamically adjusting attention parameters and enabling transfer learning across various tasks. The architecture incorporates multiple linear and affine mapping layers (Eq.2) within its feedforward network (FFN) and employs Monte Carlo Dropout (Eqs.3 and 4) instead of standard Dropout to capture richer feature representations:
$\centering y=Wx+b,\@add@centering$ (2)
where $W$ and $b$ are the weight matrix and bias vector, $x$ is the input, and $y$ is the output. During inference, multiple stochastic forward passes with dropout produce the average prediction $\hat{y}$:
@@
Designed for versatility, MMformer addresses various multivariate time series tasks, demonstrating robust adaptability and transfer learning capabilities.
@@ 3.3.2 Adaptive transferable multi-head attention
The self-attention mechanism, fundamental to the Transformer architecture, encodes each input sequence element as a high-dimensional vector [68]. For a sequence $X=(x_{1},x_{2},\dots,x_{n})$, linear transformations generate query $Q$, key $K$, and value $V$ vectors as $Q=XW_{Q}$, $K=XW_{K}$, and $V=XW_{V}$.
@@
Figure 6: Comparison between Transformer and iTransformer architectures
Fig.6 illustrates the distinct implementations of Transformer and iTransformer. Traditional Transformers aggregate multiple variables from a single time step into one token, capturing temporal dependencies but potentially hiding variable interactions. In contrast, iTransformer embeds each variable’s entire MTS into separate tokens, allowing the attention mechanism to effectively capture inter-variable correlations. This approach facilitates non-linear representations through feed-forward networks, enhancing variable-centric learning and mitigating limitations associated with mixed time tags and limited receptive fields.
Fig.7 and 8 highlight that in contrast to Transformer and iTransformer, the MMformer structure assigns independent tokens to each variable and incorporates the Adaptive Transferable Multi-Head Attention (ATMA) mechanism. ATMA integrates meta-learning to enhance multimodal data integration, capturing complex inter-variable relationships through multi-head attention. This design improves MMformer’s ability to process MTS, maintaining robustness on small datasets or those with outliers, and only requires minimal parameter tuning.
Figure 7: Comparison between Transformer and MMformer. ATMA performs forward propagation on input tasks, calculates the current batch’s loss, and accumulates it in task loss to adapt parameters to the current task. The adapted parameters are forward propagated to obtain outputs, and meta loss is calculated between outputs and targets. Meta loss is used in the outer loop to update the model’s meta parameters, while the current epoch’s validation set loss is used to update the model.
Figure 8: Comparison between iTransformer and MMformer
The ATMA framework defines a set of tasks $\mathcal{T}$, where each task $T_{i}\in\mathcal{T}$ includes a loss function $\mathcal{L}_{T_{i}}$, initial observation distribution $q(x_{1})$, transition distribution $q(x_{t+1}|x_{t},a_{t})$, and designated time step $H_{i}$. A task $T_{i}$ is randomly selected, and model parameters $\theta$ are updated to $\theta_{i}^{\prime}$ via gradient descent using loss function $\mathcal{L}_{T_{i}}$:
@@
The proposed ATMA mechanism combines self-attention with meta-learning to enhance MTS analysis. By defining a diverse task set and performing gradient updates per task, ATMA enables rapid model adaptation and improves generalization across tasks through meta-gradients. This integration extends self-attention to multi-dimensional long and short-term sequences, resulting in the MMformer model. MMformer has robustness on limited or noisy datasets while remaining user-friendly for researchers lacking tuning experience. ATMA’s innovative design offers significant potential for advancements in related domains.
@@ 4 Experiments Application and Results
This section employs the MMformer to analyze two datasets. The datasets, sourced from the China Meteorological Data Service Center under the National Meteorological Information Center [27], encompass comprehensive raw data essential for accurate forecasting and detecting trends. We benchmark MMformer against iTransformer, Transformer, and SARIMA models to demonstrate its superior performance.
This evaluation provides insights into MMformer’s effectiveness as the MTS prediction tool for environmental monitoring.
@@ 4.1 Multivariate Time Series Data Display
The proposed MMformer model’s effectiveness in predicting air quality using daily average measurements of seven indicators across 331 Chinese cities over 1,277 days (January 2018 to June 2021) and temperature patterns at 2415 monitoring sites during the summer (276 days) from 2012 to 2014. This research uses daily average values for seven air quality indicators rather than instantaneous measurements, which reduces the impact of transient phenomena on prediction outcomes. While air quality data contains ephemeral fluctuations, such as short-term pollution peaks or rapid meteorological changes induced, these phenomena are effectively smoothed in daily average aggregations, reducing extreme values’ influence on overall trends.
@@
Figure 10: Precipitation and Temperature from 2012 to 2014 June-August
Fig.10 presents a comprehensive visualization of data distribution patterns and outlier occurrences. The color gradient bar beneath each graph corresponds to temperature and precipitation values. The horizontal axis denotes temporal progression and station identification numbers, while the vertical axis represents temperature or precipitation measurements. Analysis reveals highly heterogeneous summer rainfall distribution across the three years, contrasted with consistently rising temperatures. Table 3 provides a statistical summary of the processed dataset, encompassing precipitation and temperature metrics within the climate dataset.
@@
Total Precipitation2 115 0 23 1357
2012 Precip 137 0 24 1391
2013 Precip 119 0 24 1407
2014 Precip 121 0 22 1415
Total Temperature3 301 121 243 1011
2012 Temp 328 55 242 987
2013 Temp 320 48 249 986
2014 Temp 315 46 239 949
@@
    Number of sites below mean
@@
    The unit of temperature: 0.1℃

### v2

@@ 3 Methodology
In this section, we shall first present the data preprocessing methods and then comprehensively describe the methodology, including detailed procedures and clear flowcharts. It is followed by the model’s components, detailing the innovative components. These advancements lead to substantial improvements in neural network optimization and increase the accuracy of time-series forecasting.
@@ 3.1 Data Preprocessing
  • 1.
    Step 1. Feature sorting and selection: Features are sorted and selected, aiming to identify the prediction tasks with the most relevant and influential features. This step helps to reduce noise and improve the robustness and accuracy of the model.
  • 2.
    Step 2. Missing value and outlier processing: Missing values and outliers in the dataset are treated using techniques such as linear interpolation or MTS smoothing to ensure the continuity and reliability of the data.
  • 3.
    Step 3. Data transformation and normalization: The selected features are subjected to necessary data transformation and normalization. This step aims to adjust the data to a form suitable for model processing.
  • 4.
    Step 4. Integrity check: In the final stage of preprocessing, the processed data is checked for integrity, including verification of data integrity, availability, etc., to ensure that the data quality meets the requirements of model training.
@@ 3.2 Workflow Overview
This subsection details our research methodology workflow. MMformer automates feature extraction from the dataset. During training, it employs meta-learning to learn sub-task-specific features. Subsequently, an outer loop updates shared parameters to identify and predict MTS feature patterns, as illustrated in Fig.2.
Figure 2: MMformer neural network flow chart, describing the core operating principle of the model and the operation mode of ATMA
The novelty of MMformer in workflow lies in the synergistic integration of key components, particularly in addressing the adaptability and generalization challenges inherent in complex multi-regional MTS prediction. Initially, MMformer processes the raw MTS data using a sliding-window approach to create input sequences and prediction targets, effectively generating subsets from the original data. The strategy enables the model to capture diverse temporal patterns, enhancing robustness and generalization capabilities.
A key architectural introduction involves the introduction of time stamps and dimension transposition, embedding each feature’s time series into tokens. The design choice is intended to capture inter-variable correlations more effectively. The core of MMformer’s innovation is the ATMA mechanism, which replaces the standard self-attention. ATMA uniquely integrates meta-learning (specifically, Model-Agnostic Meta-Learning, MAML). Through inner and outer optimization loops, ATMA effectively initializes and optimizes key parameters, substantially improving adaptability and generalization across various subtasks and data environments.
Furthermore, MMformer incorporates MC Dropout in the feed-forward networks (FFN), which enhances model generalization and robustness and provides uncertainty quantification, a capability not offered by traditional Dropout.
In summary, the innovation of MMformer resides in the elegant synergistic integration of deep learning (based on an encoder-only architecture) and meta-learning via the ATMA mechanism. This integration effectively addresses common challenges in complex MTS prediction, such as limited adaptability and the lack of uncertainty quantification. In contrast, it significantly reduces the reliance on manual feature engineering, improving precision and reliability.
@@ 3.3.1 Advancements in MMformer Architecture
MMformer has the advantages of the iTransformer’s architecture, enhances the self-attention mechanism, as illustrated in Fig.3. Compared with the traditional Transformer architecture and iTransformer incorporates a "Temporal LayerNorm" layer as defined in Eq.1 [18]:
where $H$ is a matrix (NN*$D$) of $N$ tokens, each with dimension $D$; $h_{n}$ is the sequence vector for token $n$, with $\text{Mean}(h_{n})$ and $\text{Var}(h_{n})$ representing its mean and variance, respectively. The purpose of expressing this operation in set form is to emphasize that standardization occurs separately for each time step (token). The set representation indicates that the same normalization procedure is applied independently to each time step $n$.
@@
Unlike conventional linear embedding approaches, MMFormer utilizes MLPs as the fundamental building blocks for both the embedding and mapping layers. The architecture utilizes within its embedding and mapping layers, processing input sequences into tokens that effectively encode both temporal and feature information. The methodology significantly improves the model’s capacity to handle diverse input types and complex temporal dynamics with greater precision. MMformer enhances MTS analysis by integrating explicit time stamp embeddings and a refined temporal positional encoding scheme. The dual approach allows the model to capture not only the relative order of data points but also their absolute temporal context.
ATMA enables the model to learn how to learn by performing quick parameter updates on a support set within an inner training loop, dynamically synchronizing parameter adjustments throughout the model, significantly enhancing its predictive capacity and generalization ability on unseen tasks. ATMA allows MMformer to leverage information from the subtasks, improving its ability to generalize to new time series forecasting tasks and significantly enhancing predictive accuracy.
Algorithm 1 MMformer Architecture: ATMA and MC Dropout
Require: Input sequences $\mathbf{X}\in\mathbb{R}^{B\times L\times C}$; Model config \$C\$; Loss function $\mathcal{L}$; Learning rates $\alpha$ (inner), $\beta$ (outer).
Ensure: Predicted sequence $\mathbf{Y}\in\mathbb{R}^{B\times L^{\prime}\times C}$.
1:Initialize model parameters $\Theta$ randomly.
2:for epoch = 1 to $N_{\text{epochs}}$ do
3: Randomly sample a mini-batch of tasks $\mathcal{T}_{i}\sim p(\mathcal{T})$ from the training tasks.
4:for each task $\mathcal{T}_{i}$ in the mini-batch do
5:  Inner Loop: Task Adaptation
6:  Sample support set $\mathcal{D}_{i}^{support}=\{(\mathbf{x}_{j},\mathbf{y}_{j})\}_{j=1}^{K}$ and query set $\mathcal{D}_{i}^{query}=\{(\mathbf{x}_{k},\mathbf{y}_{k})\}_{k=1}^{K^{\prime}}$ from $\mathcal{T}_{i}$.
7:  Initialize task-specific parameters $\Theta_{i}^{\prime}\leftarrow\Theta$.
8:  Feature Embedding:
9:   Permute $\mathbf{X}$ to $\mathbb{R}^{B\times C\times L}$ (features become tokens).
10:   Embed temporal sequences within each feature; add positional encodings.
11:   $E_{\text{enc}}\leftarrow\text{DataEmbeddingInverted}(\mathbf{X},C)$ $\triangleright$ Embeds features to $d_{\text{model}}$; attention operates across features.
12:  Encoder Stack: for $i=1$ to $N_{inner\_steps}$ do
13:   $E_{\text{enc}}\leftarrow\text{EncoderLayer}_{i}(E_{\text{enc}})$
14:   Integrates ATMA for adaptive, transferable attention: $\triangleright$ Apply Eqs.4-7
15:   Compute $\nabla_{\Theta_{i}^{\prime}}\mathcal{L}_{\mathcal{T}_{i}}(\mathbf{X},\mathbf{Y};\Theta_{i}^{\prime})$ using $\mathcal{D}_{i}^{support}$.
16:   Update parameters: $\Theta_{i}^{\prime}\leftarrow\Theta_{i}^{\prime}-\alpha\nabla_{\Theta_{i}^{\prime}}\mathcal{L}_{\mathcal{T}_{i}}$. $\triangleright$ Apply Eq.8
17:  End Inner Loop
18:  Prediction Projection:
19:   Project contextualized features ($d_{\text{model}}$) to \$L’\$ prediction length, then permute output.
20:   Apply MC Dropout. $\triangleright$ Apply Eqs.2-3
21:  Compute meta-loss $\mathcal{L}_{\text{meta}}=\mathcal{L}_{\mathcal{T}_{i}}(\mathbf{X},\mathbf{Y};\Theta_{i}^{\prime})$ using $\mathcal{D}_{i}^{query}$ and adapted parameters $\Theta_{i}^{\prime}$.
22:end for
23:Meta-Objective: Minimize meta-learning objective $\sum_{T_{i}\sim p(\mathcal{T})}\mathcal{L}_{T_{i}}\big(f_{\Theta_{i}^{\prime}}\big)$. $\triangleright$ Apply Eq.10
24:Outer Loop: Meta-Optimization
25: Update model parameters: $\Theta\leftarrow\Theta-\beta\nabla_{\Theta}\sum_{\mathcal{T}_{i}\sim p(\mathcal{T})}\mathcal{L}_{\text{meta}}$. $\triangleright$ Apply Eq.9
26:end for
27:Output: Return $\mathbf{Y}=\text{Projector}(M)$ $\triangleright$ Projects and permutes to prediction format.
To boost the model’s robustness and representational power, MC Dropout in the FFN replaces standard Dropout (see Eqs.2 and 3), which enables the capture of uncertainties in feature representations by performing multiple sampling during inference, which leads to more reliable predictions. During inference, multiple stochastic forward passes with dropout produce the average prediction $\hat{y}$:
@@
Algorithm 1 presents the operational logic of advancements in MMformer architecture, detailing its innovative structure. The pseudocode outlines data embedding, sequence encoding, ATMA, and global MC Dropout. Algorithm 1 also shows the time stamp with position and the Causal Mask mechanism of the architecture. MMformer’s architecture, with its special embedding, ATMA, and MC Dropout, constitutes a powerful novel model for MTS analysis. The innovative approach enhances predictive accuracy and strengthens the model’s generalization capabilities, offering a novel perspective and methodology for addressing complex time series problems.
@@ 3.3.2 Adaptive transferable multi-head attention
The self-attention mechanism encodes each input sequence element as a high-dimensional vector. For a sequence $X=(x_{1},x_{2},\dots,x_{n})$, linear transformations generate query $Q$, key $K$, and value $V$ vectors as $Q=XW_{Q}$, $K=XW_{K}$, and $V=XW_{V}$ [59].
@@
Figure 4: Comparison between Transformer, iTransformer, and MMformer architectures
Fig.4 illustrates the distinct implementations of Transformer, iTransformer, and MMformer. Traditional Transformers aggregate multiple variables from a single time step into one token, capturing temporal dependencies but potentially hiding variable interactions. In contrast, iTransformer embeds each variable’s entire MTS into separate tokens, allowing the attention mechanism to effectively capture inter-variable correlations. This approach facilitates non-linear representations through FFN, enhancing variable-centric learning and mitigating limitations associated with mixed time tags and limited receptive fields. However, the MMformer structure assigns independent tokens to each variable, which uses MLPs for embedding to capture the nonlinear characteristics of the data and incorporates the ATMA mechanism. The design improves MMformer’s ability to process complex multi-regional MTS, maintaining robustness on small datasets or those with outliers, and only requires minimal parameter tuning.
The ATMA module defines a set of tasks $\mathcal{T}$, where each task $T_{i}\in\mathcal{T}$ includes a loss function $\mathcal{L}_{T_{i}}$, initial observation distribution $q(x_{1})$, transition distribution $q(x_{t+1}|x_{t},a_{t})$, and designated time step $H_{i}$. A task $T_{i}$ is randomly selected, and model parameters $\theta$ are updated to $\theta_{i}^{\prime}$ via gradient descent using loss function $\mathcal{L}_{T_{i}}$:
@@
Algorithm 2 Adaptive Transferable Multi-Head Attention (ATMA)
1:Query $\mathbf{Q}$, Key $\mathbf{K}$, Value $\mathbf{V}$, Attention mask $\mathbf{M}$, Number of heads \$H\$, Model dimension \$d\$, Meta-parameters $\theta$, Learning rate $\alpha$ (inner), $\beta$ (outer)
2:Output context $\mathbf{O}$, Attention weights $\mathbf{A}$
3:Compute per-head dimension: $d_{k}=d/H$
4:Step 1: Linear Projection
[…]
8:Step 2: Meta-Learning Inner Loop (Cross-task Adaptation)
9:for each task $T_{i}$ in current meta-batch do
10:  Compute inner loss: $\mathcal{L}_{T_{i}}^{\text{train}}=\mathcal{L}(\text{ATMA}(\mathbf{Q},\mathbf{K},\mathbf{V};\theta),\penalty 10000\ \text{label}_{T_{i}})$
11:  (Meta-learning adaptation) Update task-specific parameters:
12:   $\theta^{\prime}_{T_{i}}\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{T_{i}}^{\text{train}}$ $\triangleright$ Apply Eq.8
13:end for
14:Step 3: Scaled Dot-Product Attention with MC Dropout
15:For each task $T_{i}$:
[…]
17:if $\mathbf{M}$ is not None then
18:  Mask $S$ at the specified positions by -$\infty$
19:end if
[…]
21:if training phase then
22:  Apply MC Dropout: $\mathbf{A}\leftarrow\text{MC Dropout}(\mathbf{A})$ $\triangleright$ Apply Eqs.2-3
23:end if
24:Step 4: Multi-head Context Composition and Output
[…]
26:  Concatenate $\mathbf{C}_{h}$ across all heads, reshape to $(B,L_{q},d)$
27:  Output linear layer: $\mathbf{O}=\text{Linear}_{O}(\mathbf{C})$
28:Step 5: Meta-Learning Outer Loop (Cross-task Parameter Update)
29:for each task $T_{i}$ in meta-batch do
30:  Compute validation loss: $\mathcal{L}_{T_{i}}^{\text{val}}=\mathcal{L}(\text{ATMA}(\mathbf{Q},\mathbf{K},\mathbf{V};\theta^{\prime}_{T_{i}}),\penalty 10000\ \text{label}_{T_{i}}^{\text{val}})$
31:end for
32:Aggregate meta-loss: $\mathcal{L}_{\text{meta}}=\frac{1}{|\mathcal{T}|}\sum_{T_{i}}\mathcal{L}_{T_{i}}^{\text{val}}$.
33:Define Meta-Objective: the outer loop minimizes the meta-learning objective $\sum_{T_{i}\sim p(\mathcal{T})}\mathcal{L}_{T_{i}}\big(f_{\Theta_{i}^{\prime}}\big)$. $\triangleright$ Apply Eq.10
34:Update shared parameters: $\theta\leftarrow\theta-\beta\nabla_{\theta}\mathcal{L}_{\text{meta}}$. $\triangleright$ Apply Eq.9
35:return Output $\mathbf{O}$, Attention weights $\mathbf{A}$.
@@
Algorithm 2 details the logic of the ATMA mechanism. It specifies the input parameters and outputs, and systematically outlines the process, including Linear Projection, the Meta-Learning Inner Loop, Attention with MC Dropout, Multi-head Context Composition, and the Meta-Learning Outer Loop. Algorithm 2 highlights the core innovation and comprehensively describes the entire process.
The proposed ATMA mechanism combines Probattention with meta-learning to enhance MTS analysis. By defining a diverse task set and performing gradient updates per task, ATMA enables rapid model adaptation and improves generalization across tasks through meta-gradients. This integration extends self-attention to multi-dimensional long and short-term sequences, resulting in the MMformer model. MMformer has robustness on limited or noisy datasets while remaining user-friendly for researchers lacking tuning experience. ATMA’s innovative design offers significant potential for advancements in related domains.
@@ 4 Experiments Application and Results
This section employs the MMformer to analyze two real-world datasets and general MTS datasets, PEMS. The real-world datasets, sourced from the China Meteorological Data Service Center under the National Meteorological Information Center [26], encompass comprehensive raw data. We benchmark MMformer against iTransformer, PatchTST, TimesNet, and Transformer to demonstrate its superior performance. The evaluation provides insights into MMformer’s effectiveness and generalization as the MTS prediction tool for environmental monitoring.
@@ 4.1 Multivariate Time Series Data Display
The proposed MMformer model’s effectiveness in predicting environmental change using daily average measurements of seven indicators across 331 Chinese cities over 1,277 days (January 2018 to June 2021) and temperature precipitation at 909 sites from 2018 to 2022 (1826 days). This research uses daily average values for all indicators rather than instantaneous measurements, which reduces the impact of transient phenomena on prediction outcomes. While environmental data contains ephemeral fluctuations, such as short-term pollution peaks or rapid meteorological changes, these phenomena are effectively smoothed in daily average aggregations, reducing the influence of extreme values on overall trends.
@@
Total Precipitation2 297.10 0.00 1.74 521
2018 Precip 199.88 0.00 1.76 544
2019 Precip 297.10 0.00 1.69 555
2020 Precip 205.15 0.00 1.82 547
2021 Precip 272.75 0.00 1.78 513
2022 Precip 163.34 0.00 1.64 534
Total Temperature3 37.08 -40.80 7.17 463
2018 Temp 35.56 -40.80 6.96 467
2019 Temp 36.01 -38.24 7.19 460
2020 Temp 34.08 -38.40 7.06 465
2021 Temp 36.32 -40.58 7.32 463
2022 Temp 37.08 -37.57 7.34 465
@@
    Number of sites below the mean
@@
    The unit of temperature: 1℃
Table 3 provides a statistical summary of the processed dataset, encompassing precipitation and temperature metrics within the climate dataset. Comprehensive analysis of Table 3 reveals highly heterogeneous rainfall distribution across the five years, contrasted with consistently rising temperatures.